What is the probability that a card chosen at random from a standard deck of cars will be

either a king or a heart?


1
Expert's answer
2021-05-27T10:48:09-0400

"P(King or Heart)=\frac{13+4-1}{52}=\frac{16}{52}=\frac{4}{13}."